# The inverse Kalman filter

### v5

@@ 3 Nonparametric estimation of particle interaction functions
$c^{*}_{j}(d^{*})$ $=c_{j}(d^{*},d^{*})-\bm{\Sigma}^{(u)}_{j}(d^{*})^{T}\mathbf{A}^{T}_{j}\bm{\Sigma}_{y}^{-1}\mathbf{A}_{j}\bm{\Sigma}^{(u)}_{j}(d^{*}),$ (19)
with $\bm{\Sigma}_{y}=\sum^{J}_{j=1}\mathbf{A}_{j}\bm{\Sigma}^{(u)}_{j}\mathbf{A}^{T}_{j}+\sigma^{2}_{0}\mathbf{I}_{\tilde{N}}$, $\bm{\Sigma}^{(u)}_{j}(d^{*})=[c_{j}(d^{(u)}_{1,j},d^{*}),\dots,c_{j}(d^{(u)}_{N_{j},j},d^{*})]^{T}$, and $\mathbf{r}$ representing additional system parameters.
The main computational challenge in calculating the predictive distribution in (17) lies in inverting the covariance matrix $\bm{\Sigma}_{y}$, where $\tilde{N}$ and $N_{j}$ range between $10^{5}$ and $10^{6}$ in our applications. Though various Gaussian process approximation methods have been proposed [35, 31, 8, 2, 23, 10, 3, 19], they typically approximate the parametric covariance $\bm{\Sigma}_{j}^{(u)}$ rather than $\bm{\Sigma}_{y}$, thus not directly applicable for computing the distribution in (17).
To address this, we employ the conjugate gradient (CG) algorithm [16] to accelerate the computation of the predictive distribution. Each conjugate gradient iteration needs to compute $\bm{\Sigma}_{y}\mathbf{u}$ for an $\tilde{N}$-dimensional vector ${\mathbf{u}}$. To employ the IKF, we first rearrange the input vector $\mathbf{d}^{(u)}_{j}$ into a non-decreasing input sequence $\mathbf{d}_{j}=[d_{1,j},\dots,d_{N_{j},j}]^{T}$. Denote the covariance $\mbox{var}[\mathbf{z}_{j}]=\bm{\Sigma}_{j}$ with $\mathbf{z}_{j}=[z_{j}(d_{1,j}),\dots,z_{j}(d_{N_{j},j})]^{T}$. The covariance of the observations can be written as a weighted sum of $\bm{\Sigma}_{j}$ by introducing a permutation matrix for each interaction: $\bm{\Sigma}_{y}=\sum^{J}_{j=1}\mathbf{A}_{\pi,j}\bm{\Sigma}_{j}\mathbf{A}^{T}_{\pi,j}+\sigma^{2}_{0}\mathbf{I}_{N}$, where $\mathbf{A}_{\pi,j}=\mathbf{A}_{j}\mathbf{P}^{T}_{j}$ with $\mathbf{P}_{j}$ being a permutation matrix such that $\bm{\Sigma}_{j}=\mathbf{P}_{j}\bm{\Sigma}^{(u)}_{j}\mathbf{P}^{T}_{j}$. After this reordering, the computation of $\bm{\Sigma}_{y}\mathbf{u}$ can be broken into four steps: (1) $\mathbf{u}_{j}=\mathbf{A}_{\pi,j}^{T}\mathbf{u}$, (2) $\mathbf{x}_{j}=\bm{\Sigma}_{j}\mathbf{u}_{j}$, (3) $\mathbf{\hat{u}}_{j}=\mathbf{A}_{\pi,j}\mathbf{x}_{j}$, and (4) $\sum^{J}_{j=1}\mathbf{\hat{u}}_{j}+\sigma^{2}_{0}\mathbf{u}$. Here, $\mathbf{A}_{j}$ is a sparse matrix with $N_{j}D_{y}$ non-zero entries. The IKF algorithm is used in step (2) to accelerate the most expensive computation, with all computations performed directly using terms in the KF algorithm without explicitly constructing the covariance matrix.
@@ 4.1 Evaluation criteria
For each of the $J$ latent factors, predictive performance is assessed using normalized root mean squared error (NRMSEj), the average length of the $95\%$ posterior credible interval ($L_{j}(95\%)$), and the proportion of interaction function values covered within the 95% posterior credible interval ($P_{j}(95\%)$), based on $n^{*}$ test inputs $\mathbf{d}^{*}_{j}=(d^{*}_{1,j},\dots,d^{*}_{n^{*},j})^{T}$:
@@
To account for variability due to initial particle positions, velocities, and noise, each simulation scenario is repeated $E=20$ times to compute the average performance metrics. Unless otherwise specified, parameters are estimated by cross-validation, with $80\%$ of the data used as a training set and the rest $20\%$ as the validation set. All computations are performed on a macOS Mojave system with an 8-core Intel i9 processor running at 3.60 GHz and 32 GB of RAM.
@@ 4.2 Unnormalized Vicsek model
Figure 3: (a) Computational time of calculating predictive means using IKF-CG algorithm (blue dots), CG method (purple squares), and direct computation (red triangles) for the unnormalized Vicsek model with different numbers of observations. (b, c) NRMSE of predictive means for interaction function and log-likelihood comparing direct computation (red triangles) and IKF-CG algorithm (blue dots), with differences (yellow diamonds) of order $10^{-7}$ to $10^{-5}$ and $10^{-5}$ to $10^{-4}$, respectively.
@@
where $h$ is the time step size. Particles are initialized with velocity $[v\cos(\phi_{i^{\prime}}(0)),v\sin(\phi_{i^{\prime}}(0))]^{T}$, where $\phi_{i^{\prime}}(0)$ is drawn from $\mbox{Unif}[-\pi,\pi]$ and $v={2}^{1/2}/2\approx 0.71$. Initial particle positions $\mathbf{s}_{i^{\prime}}(0)$ in (23) are uniformly sampled from $[0,{n_{p}}^{1/2}]\times[0,{n_{p}}^{1/2}]$ to keep consistent particle density across experiments. The goal is to estimate the interaction function nonparametrically, without assuming linearity. The interaction radius $r=0.5$ is estimated alongside other parameters, with the estimation results detailed in the Supplementary Material.
Figure 4: Boxplots of NRMSE (20) for estimating the latent interaction function in the unnormalized Vicsek model with $\sigma_{0}^{2}=0.1^{2}$ based on 20 experiments. Panel (a) uses the Matérn covariance with roughness parameter 2.5, and panel (b) uses the exponential covariance.
We first compare the computational cost and accuracy of IKF-CG, CG, and direct computation for calculating the predictive mean in (18) and the marginal likelihood in (16) using the covariance in (15). Simulations are conducted with $n_{p}=100$ particles and noise variance $\sigma^{2}_{0}=0.2^{2}$. Panel (a) of Fig. 3 shows that the IKF-CG approach substantially outperforms both direct computation and the CG algorithm in computational time when predicting $n^{*}=100$ test inputs with varying time lengths $n_{\tau}$, ranging from 4 to 200. Figure 3(b) compares the NRMSE of the predictive mean between the direct computation and IKF-CG method when the number of observations ranges from 400 to 2,000, corresponding to $n_{\tau}$ from 4 to 20. The two methods yield nearly identical predictive errors, with negligible differences in NRMSE. Figure 3(c) shows a comparison of the log-likelihood values computed via direct computation and IKF-CG, where the latter employs the log-determinant approximation detailed in the Supplementary Material. Both methods produce similar log-likelihood values across different sample sizes.
@@
Figure 6: (a) Predicted (blue dashed lines) and true (pink solid lines) interaction functions with $95\%$ posterior credible interval (grey shaded area) for the unnormalized Vicsek model with $n_{\tau}=900$, $T=10$, and $\sigma_{0}^{2}=0.1^{2}$, using Matérn covariance function with $\nu=2.5$. (b) Trajectories of $45$ randomly sampled particles over $50$ time frames using estimated (blue dashed lines) and true (pink solid lines) interaction functions with identical initial positions and noise samples.
Next, we evaluate the performance of the IKF-CG algorithm across 12 scenarios with varying particle numbers ($n_{p}=100,300,900$), time frames ($n_{\tau}=5,10$), and noise variances ($\sigma^{2}_{0}=0.1^{2},0.2^{2}$). The predictive performance is evaluated using $n^{*}=200$ test velocity inputs evenly spaced across a representative input domain of the interaction function in the data, $[-1,1]$, averaged over $E=20$ experiments.
Panels (a) and (b) of Figs. 4-5 show the predictive performance of particle interactions using the Matérn covariance with roughness parameter $\nu=2.5$ and the exponential covariance (Matérn with $\nu=0.5$), respectively, for noise variance $\sigma_{0}^{2}=0.1^{2}$. Results for $\sigma_{0}^{2}=0.2^{2}$ are provided in the Supplementary Material. Across all scenarios, NRMSEs remain low, with improved accuracy for larger datasets and smaller noise variances. The decrease in the average length of the $95\%$ posterior credible interval with increasing sample size, along with the relatively small interval span compared to the output range, indicates improved prediction confidence with more observations. Moreover, the coverage proportion for the $95\%$ posterior credible interval is close to the nominal $95\%$ level in all cases, validating the reliability of the uncertainty quantification. By comparing panels (a) and (b) in Figs. 4-5, we find that the Matérn covariance in (15) yields lower NRMSE and narrower $95\%$ posterior credible intervals than the exponential kernel. This improvement is due to the smoother latent process induced by (15), which is twice mean-squared differentiable, while the process with an exponential kernel is not differentiable.
@@ 4.3 A modified Vicsek model with multiple interactions
We simulate 12 scenarios, each replicated $E=20$ times, using the same number of particles, time frame, and noise level as in the unnormalized Vicsek model in Section 4.2. The predictive performance of the IKF-CG is evaluated using $n^{*}=200$ test inputs evenly spaced across the input domain of each interaction function. We focus on the model with the Matérn kernel in (15), and results for the exponential kernel are detailed in the Supplementary Material. Consistent with the results of the unnormalized Vicsek model, we find models with the Matérn kernel in (15) are more accurate due to the smooth prior imposed on the interaction function.
@@
Table 1: The predictive accuracy and uncertainty assessment by (20)-(22) for the modified Vicsek model with $\sigma_{0}^{2}=0.1^{2}$ using Matérn covariance function and roughness parameter $2.5$.
Table 1 summarizes the predictive performance for $\sigma_{0}^{2}=0.1^{2}$; results for $\sigma_{0}^{2}=0.2^{2}$ are included in the Supplementary Material. While both interaction functions exhibit relatively low NRMSEs, the second interaction function has a higher NRMSE than the first, because the repulsion at short distances creates fewer training samples with close-proximity neighbors. This discrepancy can be mitigated by increasing the number of observations. The average length of the $95\%$ posterior credible interval for the second interaction decreases as the sample size increases, and the coverage proportion remains close to the nominal 95% level across all scenarios.
Figure 7: Predictions (blue dashed curves), truth (pink solid curves), and the $95\%$ posterior credible interval (shaded region) of particle interaction functions when $n_{p}=900$, $n_{\tau}=10$, and $\sigma_{0}^{2}=0.1^{2}$.
@@ 4.4 Estimating cell-cell interactions on anisotropic substrates
We apply the IKF-CG algorithm to estimate the latent interaction function. Due to the anisotropic substrate, cellular velocities differ between horizontal and vertical directions, so we independently model the velocity of the $i^{\prime}$th cell in each direction $l$ by
@@
where $n_{p}(\tau)$ is the cell count at time $\tau$, $l=1,2$ correspond to horizontal and vertical directions, respectively, and $\epsilon_{i^{\prime},l}(\tau)\sim\mathcal{N}(0,\sigma^{2}_{0,l}(\tau))$ denotes the noise. Inspired by the modified Vicsek model, we set $d_{k,l}=v_{k,l}(\tau-1)$. To account for velocity decay caused by increasing cell confluence, we model the noise variances as $\sigma^{2}_{0,l}(\tau)=\omega_{l}\sigma^{2}_{v,l}(\tau-1)$, where $\sigma^{2}_{v,l}(\tau-1)$ is the sample velocity variance at time $(\tau-1)$, and $\omega_{l}$ is a parameter estimated from data for $l=1,2$. The neighboring set in (25) is formed as $ne_{i^{\prime},l}(\tau-1)=\{k:||\mathbf{s}_{i^{\prime}}(\tau-1)-\mathbf{s}_{k}(\tau-1)||<r_{l}\mbox{ and }\mathbf{v}_{i^{\prime}}(\tau-1)^{T}\mathbf{v}_{k}(\tau-1)>0\}$, where $r_{l}$ denotes the interaction radius of direction $l$ for $l=1,2$, excludes particles moving in opposite directions, reflecting the observed gliding and intercalating behavior of cells [37].
@@
Our nonparametric estimation of the interaction functions is compared with two models for one-step-ahead forecasts of directional velocities: the unnormalized Vicsek model introduced in Section 4.2, which uses a shared interaction radius for both velocity components, and the anisotropic Vicsek model, which uses distinct radii in each direction. As shown in Table 2, both Vicsek models have similar RMSE as the difference of interaction radii does not have a large impact for estimating interaction functions. Our model outperforms both Vicsek models in RMSE for one-step-ahead forecasts of directional velocities, despite the large inherent stochasticity in cellular motion. Moreover, our model has notably shorter average interval lengths that cover approximately $95\%$ of the held-out observations. These findings underscore the importance of the IFK-CG algorithm in enabling the use of large datasets to overcome high stochasticity and capture the underlying dynamics of cell alignment processes.

### v6

@@ 3 Nonparametric estimation of particle interaction functions
$c^{*}_{j}(d^{*})$ $=c_{j}(d^{*},d^{*})-\bm{\Sigma}^{(u)}_{j}(d^{*})^{T}\mathbf{A}^{T}_{j}\bm{\Sigma}_{y}^{-1}\mathbf{A}_{j}\bm{\Sigma}^{(u)}_{j}(d^{*}).$ (19)
Here, $\bm{\Sigma}_{y}=\sum^{J}_{j=1}\mathbf{A}_{j}\bm{\Sigma}^{(u)}_{j}\mathbf{A}^{T}_{j}+\sigma^{2}_{0}\mathbf{I}_{\tilde{N}}$, $\bm{\Sigma}^{(u)}_{j}(d^{*})=[c_{j}(d^{(u)}_{1,j},d^{*}),\dots,c_{j}(d^{(u)}_{N_{j},j},d^{*})]^{T}$ and $\mathbf{r}$ represents additional system parameters.
The main computational challenge in calculating the predictive distribution in (17) lies in inverting the covariance matrix $\bm{\Sigma}_{y}$, where $\tilde{N}$ and $N_{j}$ range between $10^{5}$ and $10^{6}$ in our applications. Though various Gaussian process approximation methods have been proposed [35, 31, 8, 2, 23, 10, 3, 20], they typically approximate the parametric covariance $\bm{\Sigma}_{j}^{(u)}$ rather than $\bm{\Sigma}_{y}$, and thus they are not directly applicable for computing the distribution in (17).
To address this, we employ the conjugate gradient (CG) algorithm [16] to accelerate the computation of the predictive distribution. Each conjugate gradient iteration needs to compute $\bm{\Sigma}_{y}\mathbf{u}$ for an $\tilde{N}-$dimensional vector $\mathbf{u}$}. To employ the IKF, we first rearrange the input vector $\mathbf{d}^{(u)}_{j}$ into a non-decreasing input sequence $\mathbf{d}_{j}=[d_{1,j},\dots,d_{N_{j},j}]^{T}$. Denote the covariance $\mbox{var}[\mathbf{z}_{j}]=\bm{\Sigma}_{j}$ with $\mathbf{z}_{j}=[z_{j}(d_{1,j}),\dots,z_{j}(d_{N_{j},j})]^{T}$. The covariance of the observations can be written as a weighted sum of $\bm{\Sigma}_{j}$ by introducing a permutation matrix for each interaction: $\bm{\Sigma}_{y}=\sum^{J}_{j=1}\mathbf{A}_{\pi,j}\bm{\Sigma}_{j}\mathbf{A}^{T}_{\pi,j}+\sigma^{2}_{0}\mathbf{I}_{N}$, where $\mathbf{A}_{\pi,j}=\mathbf{A}_{j}\mathbf{P}^{T}_{j}$ with $\mathbf{P}_{j}$ a permutation matrix such that $\bm{\Sigma}_{j}=\mathbf{P}_{j}\bm{\Sigma}^{(u)}_{j}\mathbf{P}^{T}_{j}$. After this reordering, the computation of $\bm{\Sigma}_{y}\mathbf{u}$ can be broken into four steps:
  1. (1)
    $\mathbf{u}_{j}=\mathbf{A}_{\pi,j}^{T}\mathbf{u}$,
  2. (2)
    $\mathbf{x}_{j}=\bm{\Sigma}_{j}\mathbf{u}_{j}$,
  3. (3)
    $\mathbf{\hat{u}}_{j}=\mathbf{A}_{\pi,j}\mathbf{x}_{j}$,
  4. (4)
    $\sum^{J}_{j=1}\mathbf{\hat{u}}_{j}+\sigma^{2}_{0}\mathbf{u}$.
Here, $\mathbf{A}_{j}$ is a sparse matrix with $N_{j}D_{y}$ non-zero entries. The IKF algorithm is used in step (2) to accelerate the most expensive computation, with all computations performed directly using terms in the KF algorithm without explicitly constructing the covariance matrix.
@@ 4.1 Evaluation criteria
For each of the $J$ latent factors, predictive performance is assessed using the normalized root mean squared error (NRMSEj), the average length of the $95\%$ posterior credible interval ($L_{j}(95\%)$), and the proportion of interaction function values covered within the 95% posterior credible interval ($P_{j}(95\%)$), based on $n^{*}$ test inputs $\mathbf{d}^{*}_{j}=(d^{*}_{1,j},\dots,d^{*}_{n^{*},j})^{T}$:
@@
To account for variability due to initial particle positions, velocities, and noise, each simulation scenario is repeated $E=20$ times to compute the average performance metrics. Unless otherwise specified, parameters are estimated by cross-validation, with $80\%$ of the data used as a training set and the remaining $20\%$ as the validation set. All computations are performed on a macOS Mojave system with an 8-core Intel i9 processor running at 3.60 GHz and 32 GB of RAM.
@@ 4.2 Unnormalized Vicsek model
Figure 3: (a) Computational time of calculating predictive means using the IKF-CG algorithm (blue dots), CG method (purple squares), and direct computation (red triangles) for the unnormalized Vicsek model with different numbers of observations. (b), (c) NRMSE of predictive means for the interaction function and log-likelihood computed via direct computation (red triangles) and the IKF-CG algorithm (blue dots), with differences (yellow diamonds) of orders $10^{-7}$ to $10^{-5}$ and $10^{-5}$ to $10^{-4}$, respectively.
@@
where $h$ is the time step size. Particles are initialized with velocity $v\cos(\phi_{i^{\prime}}(0)),v\sin(\phi_{i^{\prime}}(0))]^{T}$, where $\phi_{i^{\prime}}(0)$ is drawn from a uniform distribution on $\pi,\pi]$ and $v={2}^{1/2}/2\approx 0.71$. Initial particle positions $\mathbf{s}_{i^{\prime}}(0)$ in (23) are uniformly sampled from $0,{n_{p}}^{1/2}]\times[0,{n_{p}}^{1/2}]$ to keep consistent particle density across experiments. The goal is to estimate the interaction function nonparametrically, without assuming linearity. The interaction radius $r=0.5$ is estimated alongside other parameters, with the estimation results detailed in the Supplementary Material.
Figure 4: Boxplots of NRMSE in (20) for estimating the latent interaction function in the unnormalized Vicsek model with $\sigma_{0}^{2}=0.1^{2}$ based on 20 experiments. Panel (a) uses the Matérn covariance with roughness parameter 2.5, and panel (b) uses the exponential covariance.
We first compare the computational cost and accuracy of the IKF-CG, CG, and direct computation for calculating the predictive mean in (18) and the marginal likelihood in (16) using the covariance in (15). Simulations are conducted with $n_{p}=100$ particles and noise variance $\sigma^{2}_{0}=0.2^{2}$. Figure3(a) shows that the IKF-CG approach substantially outperforms both direct computation and the CG algorithm in computational time when predicting $n^{*}=100$ test inputs with varying time lengths $n_{\tau}$, ranging from 4 to 200. Figure 3(b) compares the NRMSE of the predictive mean between the direct computation and IKF-CG method when the number of observations ranges from 400 to 2,000, corresponding to $n_{\tau}$ from 4 to 20. The two methods yield nearly identical predictive errors, with negligible differences in NRMSE. Figure 3(c) shows a comparison between the log-likelihood values computed via the direct computation and the IKF-CG algorithm, where the latter employs the log-determinant approximation detailed in the Supplementary Material. Both methods produce similar log-likelihood values across different sample sizes.
@@
Figure 6: (a) Predicted (blue dashed lines) and true (pink solid lines) interaction functions with the $95\%$ posterior credible interval (grey shaded area) for the unnormalized Vicsek model with $n_{\tau}=900$, $T=10$, and $\sigma_{0}^{2}=0.1^{2}$, using Matérn covariance function with $\nu=2.5$. (b) Trajectories of $45$ randomly sampled particles over $50$ time frames using estimated (blue dashed lines) and true (pink solid lines) interaction functions with identical initial positions and noise samples.
Next, we evaluate the performance of the IKF-CG algorithm across 12 scenarios with varying particle numbers ($n_{p}=100,300,900$), time frames ($n_{\tau}=5,10$), and noise variances ($\sigma^{2}_{0}=0.1^{2},0.2^{2}$). The predictive performance is evaluated using $n^{*}=200$ test velocity inputs evenly spaced across a representative input domain of the interaction function in the data, [1,1][-1,1], averaged over $E=20$ experiments.
Panels (a) and (b) of Figs. 4-5 show the predictive performance of particle interactions using the Matérn covariance with roughness parameter $\nu=2.5$ and the exponential covariance (Matérn covariance with $\nu=0.5$), respectively, for noise variance $\sigma_{0}^{2}=0.1^{2}$. Results for $\sigma_{0}^{2}=0.2^{2}$ are provided in the Supplementary Material. Across all scenarios, NRMSEs remain low, with improved accuracy for larger datasets and smaller noise variances. The decrease in the average length of the $95\%$ posterior credible interval with increasing sample size, along with the relatively small interval span compared to the output range, indicates improved prediction confidence with more observations. Moreover, the coverage proportion for the $95\%$ posterior credible interval is close to the nominal $95\%$ level in all cases, validating the reliability of the uncertainty quantification. By comparing panels (a) and (b) in Figs. 4-5, we find that the Matérn covariance in (15) yields lower NRMSE and narrower $95\%$ posterior credible intervals than the exponential kernel. This improvement is due to the smoother latent process induced by (15), which is twice mean-squared differentiable, while the process with an exponential kernel is not differentiable.
@@ 4.3 A modified Vicsek model with multiple interactions
We simulate 12 scenarios, each replicated $E=20$ times, using the same number of particles, time frame, and noise level as in the unnormalized Vicsek model in Section 4.2. The predictive performance of the IKF-CG algorithm is evaluated using $n^{*}=200$ test inputs evenly spaced across the input domain of each interaction function. We focus on the model with the Matérn kernel in (15), with the results for the exponential kernel reported in the Supplementary Material. Consistent with the results of the unnormalized Vicsek model, we find that models with the Matérn kernel in (15) are more accurate due to the smooth prior imposed on the interaction function.
@@
Table 1: The predictive accuracy and uncertainty assessment by (20)-(22) for the modified Vicsek model with $\sigma_{0}^{2}=0.1^{2}$ using Matérn covariance function and roughness parameter $\nu_{j}=2.5$.
Table 1 summarizes the predictive performance for $\sigma_{0}^{2}=0.1^{2}$; results for $\sigma_{0}^{2}=0.2^{2}$ are reported in the Supplementary Material. While both interaction functions exhibit relatively low NRMSEs, the second interaction function has a higher NRMSE than the first, because the repulsion at short distances creates fewer training samples with close-proximity neighbors. This discrepancy can be mitigated by increasing the number of observations. The average length of the $95\%$ posterior credible interval for the second interaction decreases as the sample size increases, and the coverage proportion remains close to the nominal 95% level across all scenarios.
Figure 7: Predictions (blue dashed curves), truth (pink solid curves), and the $95\%$ posterior credible interval (shaded regions) of particle interaction functions when $n_{p}=900$, $n_{\tau}=10$, and $\sigma_{0}^{2}=0.1^{2}$.
@@ 4.4 Estimating cell-cell interactions on anisotropic substrates
We apply the IKF-CG algorithm to estimate the latent interaction function. Because of the anisotropic substrate, cellular velocities differ between horizontal and vertical directions, so we independently model the velocity of the $i^{\prime}$th cell in each direction $l$ as
@@
[…]
@@
Our nonparametric estimation of the interaction functions is compared with two models for one-step-ahead forecasts of directional velocities: the unnormalized Vicsek model introduced in Section 4.2, which uses a shared interaction radius for both velocity components, and the anisotropic Vicsek model, which uses distinct radii in each direction. As shown in Table 2, both Vicsek models have similar RMSE as the difference of interaction radii does not have a large impact for estimating interaction functions. Our model outperforms both Vicsek models in the RMSE for one-step-ahead forecasts of directional velocities, despite the large inherent stochasticity in cellular motion. Moreover, our model has notably shorter average interval lengths that cover approximately $95\%$ of the held-out observations. These findings underscore the importance of the IFK-CG algorithm in enabling the use of large datasets to overcome high stochasticity and capture the underlying dynamics of cell alignment processes.